ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 138, 2003

Aaronson, J., H. Nakada, O. Sarig and R. Solomyak Corrections to “Invariant measures and asymptotics for some skew products” by Jon Aaronson, Hitoshi Nakada, Omri Sarig and Rita Solomyak, Israel Journal of Mathematics, Vol. 128, 2002, pp. 93–134

377

Anderson, G. W. Integral Kašin splittings  139

Arroyo, J., O. J. Garay and M. Barros Closed free hyperelastic curves in the hyperbolic plane and Chen–Willmore rotational hypersurfaces

171

Barros, M. See J. Arroyo

Blagojević, P., V. Grujić and R. Živaljević Symmetric products of surfaces and the cycle index

61

Chuang, C.-L. and T.-K. Lee Algebraic derivations with constants satisfying a polynomial identity

43

Dooley, A. H. and T. Hamachi Nonsingular dynamical systems, Bratteli diagrams and Markov odometers

93

Downey, R. G., D. R. Hirschfeldt, S. Lempp and R. Solomon Computability-theoretic and proof-theoretic aspects of partial and linear orderings

271

Feng, D.-J. Lyapunov exponents for products of matrices and multifractal analysis. Part I: Positive matrices

353

Foulon, P. and B. Hasselblatt Zygmund strong foliations  157
Galvin, D. On homomorphisms from the Hamming cube to Z  189

Garay, O. J. See J. Arroyo

Grujić, V. See P. Blagojević

Guo, W., K. P. Shum and A. Skiba G-covering subgroup systems for the classes of supersoluble and nilpotent groups

125

Hamachi, T. See A. H. Dooley

Hasselblatt, B. See P. Foulon

Hirschfeldt, D. R. See R. G. Downey

Lee, T.-K. See C.-L. Chuang

Lempp, S. See R. G. Downey Mackey, M. and P. Mellon Angular derivatives on bounded

symmetric domains

291

Mann, A. and F. Posnick-Fradkin Subgroups of powerful groups  19

Mellon, P. See M. Mackey Mináč, J. and J. Swallow Galois module structure of pth-power

classes of extensions of degree p

29

Nakada, H. See J. Aaronson Pan, S.-Y. Local theta correspondence and minimal K-types of

positive depth

317

Posnick-Fradkin, F. See A. Mann Raja, M. First Borel class sets in Banach spaces and the asymptotic-

norming property

253

Sarig, O. See J. Aaronson

Schlenk, F. Symplectic embeddings of ellipsoids  215

Shum, K. P. See W. Guo Simon, R. S. Games of incomplete information, ergodic theory, and

the measurability of equilibria

73

Skiba, A. See W. Guo Solomon, R. See R. G. Downey Solomyak, R. See J. Aaronson Swallow, J. See J. Mináč

Wilson, L. E. Dimension subgroups and p-th powers in p-groups  1

Živaljević, R. See P. Blagojević